Two points in the xy plane have Cartesian coordinates (5.50, −6.00) m and (−3.50, 3.50) m


1
Expert's answer
2021-10-05T10:06:17-0400

"Distance formula =\sqrt{(x_1-x_2)^2 +(y_1-y_2)^2} \\"


"distance=\sqrt(5.5-(-3.5))^2 +(-5.5-3.5)^2 \\
=\sqrt{9^2+(-9)^2} \\
=9\sqrt {2} m"

"d=12.7279"